# Semantic Communication Meets Heterogeneous Network: Emerging Trends, Opportunities, and Challenges

### v1

@@ V-A Discrimination of Heterogeneous Model and System
Due to the heterogeneity of the semantic codec models, the traditional model aggregation approaches are not applicable to the collaborative updating of semantic codecs. Leaving heterogeneous models for local training causes incomplete information between users. Each user does not know each other’s data information, model information, and system information at the time of training. In the case of one user being slow to update due to model or system reasons, other users may exclude that user from joining the update due to considerations such as accuracy or update efficiency. This is even though that user may have more data information and potentially a higher transmission accuracy after completed updates because the other users are not informed. This prevents that user from deploying semantic communication properly due to discrimination during updating, thus necessitating a high communication load via employing conventional communication paradigam.
V-B Fairness of Heterogeneous Data
@@
Datasets of users are generally considered as data information is Non-IID. Yet, inconsistencies in the data ‘class’ not only influence updates to transmission accuracy and convergence, but also lead to an unfair SemCom after updating, since the user only gets its personalized encoder or decoder model and cannot determine the others’ decoder or encoder model. For instance, in the United Kingdom, a minority ethnic user performs collaborative semantic codec updates with other users in a community of Britons. His training information is related to his ethnicity, e.g., pictures and language. This information, however, is a minority in this community, making the updated semantic codec less accurate for him than for the local Briton community users. This creates an unfair update for the minority ethnic user, who is also involved in the update.
V-C Privacy of Personalized One-to-Many Model
@@
In the network, there is a one-to-many characteristic of semantic models, the semantic encoders and decoders that should form an integral ML model are distributed across different users. During the updating of the SemCom model, if a malicious user can get hold of the semantic encoder/decoder model of one normal user, it can not only affect the coding/decoding of that user but also cause the decoding/coding methods of other users to be deduced. This leads to leakage of transmitted information after SemCom model updating is complete, even though each user’s semantic model is individual and heterogeneous. Hence, the semantic codec model is related to the accuracy of signal transmission, users need to consider, not only transmission data privacy and system information privacy but also crucial to consider semantic codec model privacy while employing SemCom.
V-D Security of Active Defense Paradigm Changes
SemCom achieves efficient information transmission by shifting part of the communication workload to the computation workload. However, this new paradigm enables data, model, and transmission security to interact in a mutually affecting way. The utility function of attackers attacking a SemCom network changes accordingly. Furthermore, the defender’s operation and maintenance cost carries on increasing. This is because the defender has to consider not only the computational security but also the communication security. The incomplete information and complexity caused by heterogeneous scenarios exacerbate the difficulty of active defense for the defender. This coupling problem hence changes the traditional active defense analysis posture.
V-E Dynamics of Heterogeneous Network
@@
Several studies have discussed the integration of semantic communication into practical networks, which are often dynamic, e.g., vehicular networks[10] and satellite networks[14]. In such networks for semantic codec updating, there exist new users joining and veteran users leaving. The locations of users are also changed. It thus is urgently needed to address the above issues such as privacy in heterogeneous networks, while taking into account the dynamics of the user during semantic encoder updates.
V-F Enthusiasm of semantic codec Updating
@@
Facing a brand new task, users in the network associate to perform semantic codec updates to transmit spectral efficiency as well as robustness enhancements for that task. Furthermore, due to heterogeneity, each user is not equally enthusiastic in the same sense as to whether or not to transmit for this task using SemCom, i.e., whether or not to participate in the update. Nevertheless, if some users participate in the update and some refuse, it does not simply affect a single user. Because communication is mutual, users of two different communication paradigms cannot communicate. Moreover, network spectrum resources are limited, users who refuse to update take up more spectrum resources, affecting all users in the network.
(a) General aggregation
(b) Model 1
(c) Model 2
Figure 3: Updating convergence speed and transmission accuracy before/after applying our framework.
VI Conslusion
@@
This article surveyed the deployment of SemCom in heterogeneous networks, highlighting the challenges during updating semantic codecs to facilitate the widespread adoption of SemCom. Moreover, a heterogeneous semantic codec updating framework was proposed and may constitute a promising candidate for supporting SemCom in heterogeneous networks. Lastly, we also discussed some open issues, such as discrimination, fairness, privacy, network dynamics, and updating enthusiasm. We hope that the challenges and opportunities in this article will pave the way to advance SemCom for large-scale applications on the network in the future.
VII Acknowledgments
This work was supported in part by the Open Networks Ecosystem Competition Western Open Radio Access Network (O-RAN) Deployment (ONE WORD) Project.

### v2

@@ VI-A Discrimination Arising from Heterogeneous Model and System
Due to the heterogeneity of the semantic codec models, the traditional model aggregation approaches are not applicable to the collaborative updating of semantic codecs. Leaving heterogeneous models for local training causes incomplete information between users. Each user does not know each other’s data information, model information, and system information at the time of training. In the case of one user being slow to update due to model or system reasons, other users may exclude that user from joining the update due to considerations such as accuracy or update efficiency. This occurs even if the excluded user possesses richer data and could achieve higher transmission accuracy after updating, since others lack this information. This prevents this user from deploying semantic communication properly due to discrimination during updating, thus necessitating a high communication load via employing conventional communication paradigm. To mitigate this, potential solutions such as developing social-inspired anti-discrimination mechanisms that tolerate stragglers could be explored.
VI-B Fairness issues with Heterogeneous Data
@@
Datasets of users are generally considered non-IID data. Yet, inconsistencies in the data ‘class’ not only influence updates to transmission accuracy and convergence, but also lead to an unfair SemCom after updating, since the user only gets their personalized data and cannot determine the others’ data. For instance, in the United Kingdom, a minority ethnic user performs collaborative semantic codec updates with other users in a community of Britons. His training information is related to his ethnicity, e.g., pictures and language. This information, however, is a minority in this community, making the updated semantic codec less accurate for him than for the local Briton community users. This creates an unfair update for the minority ethnic user, who is also involved in the update. Potential solutions include the adoption of fairness-aware aggregation algorithms or the introduction of gradient correction mechanisms to mitigate bias against underrepresented data distributions.
VI-C Privacy Risk of Personalized M2O Model
@@
In the network, there is an M2O characteristic of semantic models, the semantic encoders and decoders that should form an integral ML model are distributed across different users. During the updating of the SemCom model, if a malicious user can get hold of the semantic encoder/decoder model of one normal user, it can not only affect the coding/decoding of that user but also cause the decoding/coding methods of other users to be deduced. This leads to leakage of transmitted information after the SemCom model updating is complete, even though each user’s semantic model is individual and heterogeneous. Hence, the semantic codec model is related to the accuracy of signal transmission. Users need to consider not only transmission data privacy and system information privacy, but also, crucially, to consider the semantic codec model privacy while employing SemCom. To enhance privacy, methods such as homomorphic encryption for model updates or differential privacy in collaborative learning could be explored.
VI-D M2O Collaborate SemCom/Inference
In the M2O SemCom paradigm, network heterogeneities critically degrade communication reliability and accuracy. It also influences the SemCom derivative paradigm, collaborative inference (CI), in which multiple devices jointly infer via semantics toward a common result. For instance, data heterogeneity challenges the fusion of disparate inputs into a coherent semantic representation, model heterogeneity introduces inconsistent feature abstractions that obscure intended meanings, and system heterogeneity corrupts transmitted semantics through varying channel conditions. These compounded effects cause the final SemCom/CI outcome to substantially deviate from the ideal, highlighting the necessity for robustness-aware M2O SemCom designs. Potential solutions could involve a graph-based semantic adapter to unify heterogeneous features.
VI-E Dynamics in Heterogeneous Network
@@
Several studies have discussed the integration of semantic communication into practical networks, which are often dynamic, e.g., vehicular networks[10] and satellite networks[13]. In such networks for semantic codec updating, there exist new users joining and veteran users leaving. The locations of users are also changed. It is thus urgently needed to address the above issues, such as privacy in heterogeneous networks, while taking into account the dynamics of the user during semantic encoder updates. Potential solutions involve designing lightweight onboarding protocols for new users and elastic aggregation strategies that adapt to network churn.
VI-F Incentive Mechanisms for Semantic Codec Updates
@@
Facing a brand new task, users in the network associate to perform semantic codec updates to transmit spectral efficiency as well as robustness enhancements for that task. Furthermore, due to heterogeneity, each user is not equally enthusiastic in the same sense as to whether or not to transmit for this task using SemCom, i.e., whether or not to participate in the update. Nevertheless, if some users participate in the update and some refuse, it does not simply affect a single user. Because communication is mutual, users of two different communication paradigms cannot communicate. Moreover, network spectrum resources are limited, and users who refuse to update. To encourage participation, incentive mechanisms such as updated credit rewards or resource allocation priority could be introduced.
VII Conclusions
@@
This article surveyed the deployment of SemCom in heterogeneous networks, highlighting the challenges during updating semantic codecs to facilitate the widespread adoption of SemCom. Moreover, a heterogeneity-aware semantic codec updating scheme was proposed and constitutes a promising candidate for supporting SemCom in heterogeneous networks. Lastly, we also discussed some open issues, such as discrimination, fairness, privacy, network dynamics, and updating enthusiasm. We hope that the challenges and opportunities in this article will pave the way to advance SemCom for large-scale applications on the network in the future.
VIII Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant 62571307, the National Key Research and Development Program of China under Grants 2022YFB2902304, the Science and Technology Commission Foundation of Shanghai under Grants 24DP1500500, and the Open Networks Ecosystem Competition Western Open Radio Access Network (O-RAN) Deployment (ONE WORD) Project. The authors would like to give special thanks to the Van Gogh Museum, Amsterdam, for granting permission to reproduce the images of Vincent van Gogh’s Sunflowers (s0031V1962) and Self-Portrait with Grey Felt Hat (s0016V1962) in this article.